Hi I need help with this
6 - 3|2x + 6| = 0

Answer by stanbon(75887)

6 - 3|2x + 6| = 0
3|2x+6| = 6
|2x+6| = 2
2x+6 = 2 or 2x+6 = -2
2x = -4 or 2x = -8
x = -2 or x = -4
===============
Cheers,
Stan H.